# Nanoscale magnetometry of a synthetic three-dimensional spin texture

### v1

@@ 4.3 Micromagnetic simulations
The multilayered SAF was modeled using Mumax3 Vansteenkiste2014 with the interlayer exchange interaction implemented as a custom field Joos2023, Salikhov2025. Four FM blocks with a thickness of 6.75 nm, separated by a 0.75 nm thick spacer, with nm3 cell size, and 8 times lateral periodic boundary conditions were used in the simulation. The structure is initially defined as a perfect square AF domain with a sharp DW (see SI). A relaxation and subsequent minimization was performed to let the magnetization evolve into the energetic minimum state. The saturation magnetization kA/m from the SQUID-VSM measurement (see supplementary figure S3) was used; other magnetic parameters were adopted from previous studies in similar multilayer structures: exchange stiffness pJ/m Salikhov2021, interlayer exchange coupling constant mJ/m2 Samad2021, perpendicular anisotropy MJ/m3 Koch2020, average grain size 32 nm with a normally distributed anisotropy direction variation among the grains Salikhov2022. Second-order anisotropy terms and interfacial DMI were not considered in the model.
This work was supported by an Emmy Noether grant from the Deutsche Forschungsgemeinschaft (DFG), project No. 504973613. R.J.P.R and A.S acknowledge financial support by the DFG under the Germany’s Excellence Strategy through the Würzburg-Dresden Cluster of Excelence on Complexity and Topology Quantum Matter-ct.qmat (EXC 2174, project-id 390858490). Furthermore, this work was supported by the DFG through the project No. 514946929 at Chemnitz Univeristy of Technology.
See pages - of Supp_Info.pdf

### v2

@@ 4.3 Micromagnetic simulations
The multilayered SAF was modeled using Mumax3 52 with the interlayer exchange interaction implemented as a custom field 25, 45. Four FM blocks with a thickness of 6.75 nm, separated by a 0.75 nm thick spacer, with nm3 cell size, and 8 times lateral periodic boundary conditions were used in the simulation. The simulation volume includes large lateral dimensions (2.048 m × 2.048 m) to match the experimental field of view and extends 75 nm above the sample surface to capture variations of the stray field within the estimated range for the NV-sample distance. To evaluate possible mesh-size effects in the description of the spin texture, a lateral mesh resolution of 2 nm was also tested, which resolves additional small-scale variations but does not significantly change the overall magnetization configuration. This high-resolution discretization of the mesh, however, increases both memory consumption and computation time. Therefore, the lateral cell size of 4 nm was chosen to balance the large simulation volume with manageable computation time.
The structure is initially defined as a perfect square AF domain with a sharp DW (see SI Figure S11). A relaxation and subsequent minimization was performed to let the magnetization evolve into the energetic minimum state. The saturation magnetization kA/m from the SQUID-VSM measurement (see SI Figure S3) was used; other magnetic parameters were adopted from previous studies in similar multilayer structures: exchange stiffness pJ/m 44, interlayer exchange coupling constant mJ/m2 46, perpendicular anisotropy MJ/m3 28, average grain size 32 nm with a normally distributed anisotropy direction variation among the grains 43. Second-order anisotropy terms and interfacial DMI were not considered in the model.
5 Supporting Information
Characterization of NV probes; determination of the NV-sample distance; MFM images and magnetic hysteresis loop; additional NV-PL maps and ODMR measurements; estimation of contrast in NV-PL maps exhibiting MFM-like contrast; iso-B images; vector magnetic field reconstruction; close-up view of the ODMR contrast on AF domains; micromagnetic simulation codes and magnetization simulations; T1 measurements; NV saturation curve and NV-PL maps at different laser powers.
This work was supported by an Emmy Noether grant from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation), project No. 504973613. R.J.P.R. and A.S. acknowledge financial support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through the Würzburg–Dresden Cluster of Excellence ctd.qmat – Complexity, Topology and Dynamics in Quantum Matter (EXC 2147, project-id 390858490). Furthermore, this work was supported by the DFG through the project No. 514946929 at Chemnitz University of Technology.